OpenStudy (cheesymcdouble):

the midpoint of cd is e(-1,0). one endpoint is c(5,2). what are the coordinates of the other endpoint?

OpenStudy (freckles):

you want the average of the x values of the endpoints to equal the x value of the midpoint you want the average of the y values of the endpoints to equal the y value of the midpoint remember the average of two numbers is the sum of those numbers divided by 2

OpenStudy (freckles):

so you have the following two equations to solve: \[\frac{5+x}{2}=-1 \text{ and } \frac{2+y}{2}=0\]

OpenStudy (freckles):

where (x,y) is assumed to be the other endpoint

OpenStudy (cheesymcdouble):

ok thanks that really helps a lot

OpenStudy (freckles):

np :)